# Pierce the Mists, Greet the Sky: Decipher Knowledge Overshadowing via Knowledge Circuit Analysis

### v2

@@ 5 Conclusion
This paper investigates hallucinations in LLMs caused by knowledge overshadowing, and introduces PhantomCircuit, a novel knowledge circuit-based analysis framework. PhantomCircuit first analyzes the training dynamics of overshadowing, finding that dominant knowledge popularity, model size, and dataset size critically shape the onset, duration, and recovery of overshadowing. Apart from that, the persistent overshadowing in pretrained models stems from inadequately optimized subordinate knowledge loss. By analyzing knowledge circuits, we find that changes in critical attention heads’ focus on subordinate knowledge directly correlate with the recovery or onset of overshadowing. Finally, optimizing these knowledge circuits presents a promising strategy for mitigating knowledge overshadowing.
@@
    The dynamic analysis of knowledge circuits throughout training is computationally intensive, potentially hindering scalability to very large models or extensive training. We aim to develop more computationally efficient techniques for approximating circuit evolution, such as checkpoint-based analysis or lightweight probing.
@@
    This study concentrates on a specific type of knowledge overshadowing, leaving more complex or subtle interference patterns unaddressed. Future work will broaden PhantomCircuit’s scope to investigate a wider range of overshadowing phenomena, including those in multi-hop reasoning.
@@
    Future efforts will focus on evolving our instance-specific circuit optimizations into a generalized mitigation toolkit, supported by a comprehensive evaluation framework. Key improvements will target the precision of automated overshadowed knowledge identification and the broader efficacy of circuit-based interventions. Ultimately, we aim to develop PhantomCircuit as a robust platform for both in-depth analysis and effective, generalizable overshadowing mitigation.
@@ A.1 Large Language Models
First proposed by Brown et al. (2020), Transformer-based auto-regressive LLMs have demonstrated strong performance across a variety of NLP tasks, including question answering Yue (2025), in-context learning Dong et al. (2022), and analogical reasoning Webb et al. (2023). Pretrained on large-scale text corpora, LLMs have acquired extensive real-world knowledge from web sources. As a result, models such as InternLM2.5 Cai et al. (2024), Qwen2.5 Team (2024), and LLaMA3.3 Grattafiori et al. (2024) have shown excellent performance on world knowledge benchmarks Suzgun et al. (2022). Therefore, over the past year, LLMs have demonstrated remarkable capabilities in understanding-related tasks across various fields Kim et al. (2024); Nam et al. (2024); Yan et al. (2024c); Yan and Lee (2024); Yan et al. (2024b); Dong et al. (2025); Su et al. (2025).
@@ B.1 Circuit Construction
We adapted the optimized circuit construction method provided by Yao et al. (2024). The process begins by representing the LLM as a directed acyclic graph (DAG), $G=(V,E)$, where $V$ encompasses input embeddings, attention heads, MLP layers, and output logits, and $E$ represents the information flow between these components. The goal is to identify a subgraph $C\subseteq G$ that is critical for recognize the key component of a given input prompt, particularly in knowledge overshadowing, is $X_{dom}$ and $X_{sub}$, the difference between $P_{dom}$ and $P_{sub}$.
The adapted construction method is similar to edge attribution patching(EAP) Conmy et al. (2023), which involves:
@@
    where $Y_{target}$ is ideally $Y_{sub}$. The expectation $\mathbb{E}$ is taken over instances of $P_{sub}$ in our evaluation set $Z_{sub}$. In practice, methods like Integrated Gradients (IG) are often used to refine this attribution by integrating gradients along a path from a baseline (corrupted) input to the actual (clean) input.
@@ B.2 Automated Component Identification for Recovery
$S_{R\textnormal{-}PMI^{-}}(P_{sub},P^{\prime}_{sub})=\sum\min(R\textnormal{-}PMI(y_{i}),0).$
The $X^{\prime}_{sub}$ that yields the minimum (most negative) $S_{R\textnormal{-}PMI^{-}}$ is identified as the primary overshadowed component, $X_{sub}^{*}$. This selection is based on the rationale that removing the true $X_{sub}^{*}$ most strongly exposes the model’s bias towards outputs favored by the dominant knowledge pattern.
Identifying Target Subordinate Output $Y_{sub}$. The intended subordinate output $Y_{sub}$ is identified by assessing which token from $V_{top}(P_{sub})$ (the top-k candidates for the original prompt $P_{sub}=(X_{bg},X_{sub})$) exhibits the most significant improvement when the overshadowing influence of background knowledge ($X_{bg}$) or other non-subordinate components is mitigated. Specifically, we generate contrastive prompts $P^{\prime}_{sub}$ by masking or altering components of $X_{bg}$ (or other identified non-subordinate elements that contribute to the $X_{bg}\leftrightarrow Y_{dom}$ association) within $P_{sub}$. $Y_{sub}$ is then the token $y_{i}\in V_{top}(P_{sub})$ that shows the most substantial rank improvement (or largest increase in log probability) in these modified prompts $P^{\prime}_{sub}$ compared to its rank in the original $P_{sub}$. This rank elevation signifies the “unmasking” of the true subordinate answer as the dominant, overshadowing associations are weakened.
@@
[…]
@@
{tblr}
[…]
@@
[…]
@@
[…]
@@
[…]
@@ C.1 Detailed Synthetic Dataset Construction
C.2 Finetuning dataset
For the finetuning dataset, we utilized the Qwen-Long API to generate instances of virtual knowledge. This generated data subsequently underwent manual review to identify and remove any instances that are overly repetitive or semantically too similar, ensuring a degree of diversity, resulted in D = 1M.
A key distinction from the synthetic dataset construction is that we did not strictly control token lengths for each component in this dataset. Instead of randomly sampled token IDs, the finetuning dataset consists of actual linguistic statements that, while syntactically and semantically coherent, represent virtual (i.e., fabricated but plausible) knowledge. The underlying pattern of dominant and subordinate knowledge construction, however, mirrors that of the synthetic dataset.
@@ C.2 Finetuning dataset
Table 3: Illustrative examples from the Finetuning Dataset (P=5). Each data entry is a row, with fine lines separating entries within a group.
{tblr}
[…]
@@
[…]
@@
[…]
@@
(a) The $\mathcal{RO}$ during training phase of the different finetuning dataset size (D).
(b) The co-evolution of $\mathcal{LP}$ and $\mathcal{RO}$ during training phase on finetuning dataset.
Figure 6: The dynamics analysis of knowledge overshadowing in finetuning dataset.
Appendix D More Dynamics Analysis on Finetuning Dataset
@@
In addition to validating the efficacy of our circuit-based overshadowing recovery method, the finetuning dataset serves a dual purpose. We also leverage it to empirically verify our conclusions regarding the training dynamics of knowledge overshadowing, specifically concerning the impact of Dataset Size (D). Consistent with our dynamic analysis findings, we investigate whether a larger D indeed correlates with a slower recovery rate from knowledge overshadowing and a prolonged duration of the hallucination effect. To this end, we conduct experiments on the finetuning dataset by fixing Knowledge Popularity at P=5 and Model Size at M=70M, while varying D across values of 0.1M, 0.5M, and 1M tokens. The results, as depicted in Figure 6(a), corroborate this relationship. Furthermore, under the specific configuration of P=5, M=70M, and D=0.5M on the finetuning dataset, we re-examine the interplay between the loss proportion of subordinate knowledge ($\mathcal{LP}$) and the relative overshadowing rate ($\mathcal{RO}$). As shown in Figure 6(b), the observations again support the hypothesis that insufficient optimization of subordinate knowledge contributes to the persistence of knowledge overshadowing.
@@
It is noteworthy that distinct behaviors are observed when comparing the finetuning dataset to the synthetic dataset. Firstly, the recovery from overshadowing on the finetuning dataset is generally slower than on the synthetic dataset for same D. This can be attributed to the richer semantic relationships and greater complexity inherent in the natural language of the finetuning data, which presents a more challenging learning task.
@@
Secondly, we observe that the finetuning dataset exhibits a minimal or absent onset phase for knowledge overshadowing, where $\mathcal{RO}$ typically rise. This is because finetuning commences from a pretrained model, which has already moved beyond the initial epochs of chaotic, random predictions. Consequently, the model can very rapidly generalize strong association patterns present in the finetuning data. Moreover, the diverse and varied forms of data within the finetuning set may act akin to a beneficial noise signal, prompting the model to pay closer attention to distinguishing features and differences. This inherent data diversity can help preemptively mitigate or even eliminate the early onset stage of knowledge overshadowing that might otherwise be observed.

### v4

@@ 5 Conclusion
This paper investigates hallucinations in LLMs caused by knowledge overshadowing, and introduces PhantomCircuit, a novel knowledge circuit-based analysis framework. PhantomCircuit first analyzes the training dynamics of overshadowing, finding that dominant knowledge popularity, model size, and dataset size critically shape the onset, duration, and recovery of overshadowing. Apart from that, the persistent overshadowing in pre-trained models stems from inadequately optimized subordinate knowledge loss. By analyzing knowledge circuits, we find that changes in critical attention heads’ focus on subordinate knowledge directly correlate with the recovery or onset of overshadowing. Finally, optimizing these knowledge circuits presents a promising strategy for mitigating knowledge overshadowing.
@@
    Despite of the experiments on gpt2-medium and natural data case (Appendix E), this study’s findings primarily rely on the Pythia model suite and synthetic/fine-tuning datasets, which are chosen to ensure architectural consistency and data purity for controlled analysis. It remains to be validated whether these conclusions can be directly generalized to other model architectures and more complex, noisy real-world data.
@@
    The current work concentrates on direct knowledge interference arising from imbalanced knowledge popularity. The applicability of PhantomCircuit to practical tasks like multi-hop reasoning and code generation, remains a key area for future exploration.
@@
    Performing dynamic analysis of knowledge circuits throughout the entire training process is computationally intensive, which could potentially hinder its scalability to very large models or extensive training runs. Future work aims to develop more computationally efficient techniques, such as more efficient circuit extraction methods like Information Flow Routes (IFR), to mitigate this challenge.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No.62506318); Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); CAAI-Ant Group Research Fund; Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
@@ A.1 Large Language Models
First proposed by Brown et al. (2020), Transformer-based auto-regressive LLMs have demonstrated strong performance across a variety of NLP tasks, including question answering Yue (2025), in-context learning Dong et al. (2022), and analogical reasoning Webb et al. (2023). pre-trained on large-scale text corpora, LLMs have acquired extensive real-world knowledge from web sources. As a result, models such as InternLM2.5 Cai et al. (2024), Qwen2.5 Team (2024), and LLaMA3.3 Grattafiori et al. (2024) have shown excellent performance on world knowledge benchmarks Suzgun et al. (2022). Therefore, over the past year, LLMs have demonstrated remarkable capabilities in understanding-related tasks across various fields Kim et al. (2024); Nam et al. (2024); Yan et al. (2024c); Yan and Lee (2024); Yan et al. (2024b); Dong et al. (2025); Su et al. (2025).
@@ B.1 Circuit Construction
We adapted the optimized circuit construction method provided by Yao et al. (2024). The process begins by representing the LLM as a directed acyclic graph (DAG), $G=(V,E)$, where $V$ encompasses input embeddings, attention heads, MLP layers, and output logits, and $E$ represents the information flow between these components. Our goal is to identify a subgraph $C$ that is critical for recognizing the key component of a given input prompt. In the context of knowledge overshadowing, this component is $\{X_{dom},X_{sub}\}$, which represents the difference between $P_{dom}$ and $P_{sub}$.
The adapted construction method is similar to Edge Attribution Patching with Integrated Gradients (EAP-IG) Hanna et al. (2024), which involves:
@@
    where $Y_{target}$ is ideally $Y_{sub}$. The expectation $\mathbb{E}$ is taken over instances of $P_{sub}$ in evaluation set. In practice, methods like Integrated Gradients (IG) are often used to refine this attribution by integrating gradients along a path from a baseline (corrupted) input to the actual (clean) input.
@@
  5. 5.
    Threshold $\tau$ Optimization: By constructing a circuit for a given edge pruning threshold, $\tau$, and then evaluating its performance metric, $M$, we can establish the relationship between $\tau$ and $\mathcal{M}$. Therefore, optimizing the circuit is equivalent to finding an optimal threshold, $\tau_{opt}$, that maximizes the performance metric $\mathcal{M}$. In practice, however, we directly optimize the number of edges $n$ rather than the threshold $\tau$. We divide this optimization process into two stages. In the first stage, we start from an initial edge count and incrementally increase the number of edges at uniform intervals to map out the relationship between $\mathcal{M}$ and $n$. We find that this relationship is multi-modal, as shown in Figure 8. In the second stage, we employ the golden-section search algorithm within the most promising interval identified previously to precisely locate the optimal number of edges, $n_{opt}$. This process yields the optimized circuit, $C_{opt}(n_{opt})$.
@@ B.2 Automated Component Identification for Recovery
$S_{R\textnormal{-}PMI}(P_{sub},P^{\prime}_{sub})=\sum\min(R\textnormal{-}PMI(y_{i}),0).$
The $X^{\prime}_{sub}$ that yields the minimum (most negative) $S_{R\textnormal{-}PMI}$ is identified as the primary overshadowed component, $X_{sub}^{*}$. This selection is based on the rationale that removing the true $X_{sub}^{*}$ most strongly exposes the model’s bias towards outputs favored by the dominant knowledge pattern.
In practical implementation, we enhance the identification of $X_{sub}^{*}$ by calculating a weighted $S_{R\textnormal{-}PMI}$. To do this, we weight each token’s $R-PMI(y_{i})$ value by the variance of its log probability change, $var(y_{i})$, observed when masking different candidates of $X_{sub}$. This results in $S_{R\textnormal{-}PMI}(P_{sub},P^{\prime}_{sub})_{weighted}=\sum\min(R\textnormal{-}PMI(y_{i})\cdot Var(y_{i}),0)$, which amplifies the effect from the $Y_{s}ub$ token which is most sensitive to the masking of $X^{*}_{sub}$, enabling a more robust identification.
Identifying Target Subordinate Output $Y_{sub}$. The intended subordinate output $Y_{sub}$ is identified by assessing which token from $V_{top}(P_{sub})$ (the top-k candidates for the original prompt $P_{sub}=(X_{bg},X_{sub})$) exhibits the most significant improvement when the overshadowing influence of background knowledge ($X_{bg}$) or other non-subordinate components is mitigated. Specifically, we generate contrastive prompts $P^{\prime}_{sub}$ by masking or altering components of $X_{bg}$ within $P_{sub}$. $Y_{sub}$ is then the token $y_{i}\in V_{top}(P_{sub})$ that shows the most substantial rank improvement in these modified prompts $P^{\prime}_{sub}$ compared to its rank in the original $P_{sub}$. Furthermore, to properly value the rank improvements of initially high-ranked tokens like $Y_{dom}$ and $Y_{sub}$, which have limited room for elevation, we introduce a weight to measure the original rank’s impact. This yields the final weighted rank elevation metric for every predicted top-$k$ next-tokens: $\text{Weighted Rank Elevation}=(k-\text{rank}_{\text{original}})\times\text{Avg. Elevation of Rank}.$
This approach preserves the effect of rank improvements for top-ranked tokens. And the rank elevation signifies that the mitigation of impact from $X_{dom}\leftrightarrow X_{bg}$ make overshadowing association weaken, which improves the rank of $Y_{sub}$.
@@
With $X_{bg}$ (background knowledge), $X_{sub}$ (identified subordinate component), the expected $Y_{sub}$, and the interfering $Y_{dom}$ established, we prepare the paired inputs required for knowledge circuit construction. The clean input is the original subordinate prompt $P_{sub}=(X_{bg},X_{sub})$, for which the desired output is $Y_{sub}$. To create the corrupt input $P_{dom}$, which is designed to elicit the overshadowing effect and output $Y_{dom}$, we maintain the background knowledge $X_{bg}$ but replace the subordinate component $X_{sub}$ with a generic placeholder token, such as ’something’. Thus, $P_{dom}=(X_{bg},\text{``something''})$. This specific formulation of $P_{dom}$ ensures that while the input structure is similar to $P_{sub}$, the absence of $X_{sub}$ allows the strong $X_{bg}\leftrightarrow Y_{dom}$ association to dominate, leading to the incorrect prediction $Y_{dom}$. These paired inputs, $P_{sub}$ and $P_{dom}$, then serve as the foundation for the activation difference calculations in our circuit analysis.
@@
Case $P_{sub}$ with $\{X_{sub}\}$ $\mathcal{M}$ indicator (logits difference) $Y_{dom}\;\&\;Y_{sub}$ Full Model Top 5 Prediction Circuit Top 5 Prediction
Case 1 Analysis of the Chrono-Filter device efficiency for temporal sorting shows outcome {filtration overload} Original model:-1.283 & Circuit: 0.764 |Time| & |Tem| Rank 0: Logit: 16.18 Prob: 32.18% Token: |Tem| Rank 0: Logit: 21.03 Prob: 73.70% Token: |Time|
@@
Case 2 Constructing psionic wave emitters necessitates precise tuning involving specialized harmonic {feedback loop} Original model: -0.745& Circuit: 2.490 |Wave| & |E| Rank 0: Logit: 17.64 Prob: 37.18% Token: |Wave| Rank 0: Logit: 17.96 Prob: 46.34% Token: |E|
@@
Case 3 Analyzing Ectoplasmic Conduit energy transfer efficiency through degrading {structure reinforcement reveals} Original model: -3.019 & Circuit: 0.523 |Transfer| & |Emit| Rank 0: Logit: 49.08 Prob: 49.26% Token: |Transfer| Rank 0: Logit: 44.54 Prob: 47.18% Token: |Emit|
@@
Case 4 Shard Relic residual energy output response to sudden energy {conduit field spikes} shows Original model: -2.623 & Circuit: 3.202 |Output| & |St| Rank 0: Logit: 40.96 Prob: 91.94% Token: |Output| Rank 0: Logit: 34.61 Prob: 51.02% Token: |St|
@@ Appendix C Dataset Details
(a) The $\mathcal{RO}$ during training phase of the different fine-tuning dataset size (D).
(b) The co-evolution of $\mathcal{LP}$ and $\mathcal{RO}$ during training phase on fine-tuning dataset.
Figure 7: The dynamics analysis of knowledge overshadowing in fine-tuning dataset.
@@ C.1 Detailed Synthetic Dataset Construction
C.2 fine-tuning dataset
For the fine-tuning dataset, we utilized the Qwen-Long API to generate instances of virtual knowledge. This generated data subsequently underwent manual review to identify and remove any instances that are overly repetitive or semantically too similar, ensuring a degree of diversity, resulted in D = 1M.
A key distinction from the synthetic dataset construction is that we did not strictly control token lengths for each component in this dataset. Instead of randomly sampled token IDs, the fine-tuning dataset consists of actual linguistic statements that, while syntactically and semantically coherent, represent virtual (i.e., fabricated but plausible) knowledge. The underlying pattern of dominant and subordinate knowledge construction, however, mirrors that of the synthetic dataset.
@@
Table 4: Illustrative examples from the fine-tuning Dataset (P=5). Each data entry is a row, with fine lines separating entries within a group.
Group $X_{bg}$ $X_{dom}$ $Y_{dom}$ $X_{sub}$ $Y_{sub}$
Group 1 Analysis of the Starfire Crystal Engine reveals primary energy output peak resonance Pure Nova
@@
Group 2 Observation of the Chronos Orb’s temporal distortion signature indicates full moonlight Stable Chrono-displacement
@@
Group 3 Study of the Xylosian biology reveals primary nutrient absorption method consuming light Molecular Assimilation
@@
(a) gpt2-medium
(b) pythia-410m
Figure 8: The relationship between circuit performance, high-attention heads proportion, average attention score on $X_{sub}$ and the number of edges in gpt2-medium and pythia-410m.
Appendix D More Dynamics Analysis on fine-tuning Dataset
@@
In addition to validating the efficacy of our circuit-based overshadowing recovery method, the fine-tuning dataset serves a dual purpose. We also leverage it to empirically verify our conclusions regarding the training dynamics of knowledge overshadowing, specifically concerning the impact of Dataset Size (D). Consistent with our dynamic analysis findings, we investigate whether a larger D indeed correlates with a slower recovery rate from knowledge overshadowing and a prolonged duration of the hallucination effect. To this end, we conduct experiments on the fine-tuning dataset by fixing Knowledge Popularity at P=5 and Model Size at M=70M, while varying D across values of 0.1M, 0.5M, and 1M tokens. The results, as depicted in Figure 7(a), corroborate this relationship. Furthermore, under the specific configuration of P=5, M=70M, and D=0.5M on the fine-tuning dataset, we re-examine the interplay between the loss proportion of subordinate knowledge ($\mathcal{LP}$) and the relative overshadowing rate ($\mathcal{RO}$). As shown in Figure 7(b), the observations again support the hypothesis that insufficient optimization of subordinate knowledge contributes to the persistence of knowledge overshadowing.
@@
It is noteworthy that distinct behaviors are observed when comparing the fine-tuning dataset to the synthetic dataset. Firstly, the recovery from overshadowing on the fine-tuning dataset is generally slower than on the synthetic dataset for same D. This can be attributed to the richer semantic relationships and greater complexity inherent in the natural language of the fine-tuning data, which presents a more challenging learning task.
@@
Secondly, we observe that the fine-tuning dataset exhibits a minimal or absent onset phase for knowledge overshadowing, where $\mathcal{RO}$ typically rise. This is because fine-tuning commences from a pre-trained model, which has already moved beyond the initial epochs of chaotic, random predictions. Consequently, the model can very rapidly generalize strong association patterns present in the fine-tuning data. Moreover, the diverse and varied forms of data within the fine-tuning set may act akin to a beneficial noise signal, prompting the model to pay closer attention to distinguishing features and differences. This inherent data diversity can help preemptively mitigate or even eliminate the early onset stage of knowledge overshadowing that might otherwise be observed.
Appendix E More Analysis on Natural Language Data
To further validate the generalization of our findings, we conduct experiments on the pre-trained Pythia-410m and gpt2-medium models using natural language prompts. The paired inputs for this test are: $P_{dom}$: "The name of this North Korean politician is", $Y_{dom}$: "Kim Jong Un", $X_{dom}$: "politician"} and $P_{sub}$: "The name of this North Korean singer is", $Y_{sub}$: "Hyon Song-wol", $X_{sub}$: "singer"}. Since it is hard to reproduce the pre-training process of these models, we instead used the number of edges in the circuit as the independent variable. We investigated the relationship among three metrics: circuit performance, the proportion of high-attention heads, and the average attention score on $X_{sub}$.
The results are shown in Figure 8, which lead to the following conclusions:
  • The model’s overall attention to the different parts $\{X_{dom},X_{sub}\}$ increases as the number of edges is reduced through optimization. Correspondingly, the model’s performance improves with this rise in attention. This validates our conclusion that higher attention on $\{X_{dom},X_{sub}\}$ is key to mitigating knowledge overshadowing.
  • The proportion of high-attention heads also increases during edge optimization. This supports the conclusion that retaining these high-attention heads plays a crucial role in directing the model’s focus toward $\{X_{dom},X_{sub}\}$.
  • The model’s performance exhibits a notable multi-modal relationship with the number of edges, which makes the optimization of the circuit’s edge count a more complex task.
Additionally, our preliminary exploration suggests that knowledge overshadowing can also occur in multi-hop reasoning. When the pre-trained OLMo-7B-Instruct model is given the prompt: "The old country of this famous scientist who is not famous for theory of relativity is", it outputs "Germany". Furthermore, with the prompt: "The name of this scientist who is not famous for theory of relativity is", the model outputs "Albert". We can infer from this that the model still defaults to "Albert Einstein" as the basis for its multi-hop reasoning.